# Exploiting Symmetry in Dynamics for Model-Based Reinforcement Learning with Asymmetric Rewards

### v2

@@ Abstract
Recent work in reinforcement learning has leveraged symmetries in the model to improve sample efficiency in training a policy. A commonly used simplifying assumption is that the dynamics and reward both exhibit the same symmetry; however, in many real-world environments, the dynamical model exhibits symmetry independent of the reward model. In this paper we assume only the dynamics exhibit symmetry, extending the scope of problems in reinforcement learning and learning in control theory where symmetry techniques can be applied. We use Cartan’s moving frame method to introduce a technique for learning dynamics which, by construction, exhibit specified symmetries. Numerical experiments demonstrate that the proposed method learns a more accurate dynamical model.
@@ I Introduction
One idea gaining recent attention in RL methods is to integrate symmetries into model-based information. Using knowledge of symmetries, one can extrapolate observed behavior to other scenarios. Recent work has leveraged symmetries via equivariant networks [vanderpol2021mdp, wang_mathrmso2-equivariant_2022], which enforce certain transformation relationships on the inputs and outputs, and via data augmentation [weissenbacher2022koopman], where artificial data is created based on observed data and known transformations for training. However, a limitation of equivariant networks is that they must be designed for the particular group of symmetries to which they are equivariant. Unlike the above methods, which assume that the symmetries are known apriori, reference [weissenbacher2022koopman] learns symmetries from data and augments the training data using these symmetries.
@@
A common assumption when using symmetry in RL and optimal control is that both the dynamics and reward (or cost) function exhibit the same symmetries, which results in the optimal policy (or control law) likewise being symmetric [ravindran_model_2002, maidens_dynamic_programming]. The symmetric policy structure is exploited for Markov Decision Processes (MDPs) and RL in [vanderpol2021mdp, wang_mathrmso2-equivariant_2022, ravindran_model_2002], [zinkevich_symmetry_2001], and for dynamic programming in [maidens_dynamic_programming]. In many scenarios, however, symmetries in the dynamical model are not necessarily applicable to the reward model. For example, multiple tasks might be executed in the same environment, where the same dynamical symmetries exist, while the reward model changes from task to task. Exploiting symmetries only in the dynamical model widens the scope of RL problems to which symmetry can be applied.
@@
This paper is organized as follows: in Section II, we provide background information on symmetries, and then use Cartan’s method of moving frames to relate a symmetrical dynamical model to a function operating on an input space of reduced dimension. In Section III, we use the result of Section II to learn a dynamical model which, by construction, satisfies specified symmetries. In Section IV we apply symmetry reduction to learn dynamical models for two examples.
@@ II Symmetries in Model Dynamics
We represent symmetries of (1) with a $r$-dimensional Lie group $G$ which is a Lie transformation group of both $\mathcal{X}$ and $\mathcal{U}$. Assume $r\leq n$. For notation, we consider the maps $\phi_{g}:\mathcal{X}\to\mathcal{X}$ defined by $x\mapsto g\cdot x$ and $\psi_{g}:\mathcal{U}\to\mathcal{U}$ defined by $u\mapsto g\cdot u$. Since the group actions are smooth, both of these maps are smooth for all $g\in G$. Further, they have smooth inverses $\phi^{-1}_{g}=\phi_{g^{-1}}$ and $\psi^{-1}_{g}=\psi_{g^{-1}}$.
@@ II-A Cartan’s Moving Frame
Consider the Lie group $G$ and its action on $\mathcal{X}$ as before. Assume $\mathcal{X}$ can be split into $\mathcal{X}^{a}$ and $\mathcal{X}^{b}$ with $r$ and $n-r$ dimensions respectively such that, for some $x_{0}\in\mathcal{X}$, the map from $g\in G$ to the projection of $g\cdot x_{0}$ onto $\mathcal{X}^{a}$ is invertible. Pick $c\in\mathcal{X}^{a}$ in the range of this map. For all $g\in G$, let $\phi_{g}^{a}:\mathcal{X}\to\mathcal{X}^{a}$ and $\phi_{g}^{b}:\mathcal{X}\to\mathcal{X}^{b}$ denote the projections of $\phi_{g}$ onto $\mathcal{X}^{a}$ and $\mathcal{X}^{b}$ respectively. Assume that, for all $x\in\mathcal{X}$, there exists a unique $g\in G$ such that $\phi_{g}^{a}(x)=c$. The set $\mathcal{C}=\{x|\phi_{e}^{a}(x)=c\}$ is an $n-r$ dimensional submanifold of $\mathcal{X}$ called a cross-section. We can then define the unique map $\gamma:\mathcal{X}\to G$, such that $\phi_{\gamma(x)}(x)\in\mathcal{C}$. This $\gamma$ is called a moving frame. Using $\gamma$, we define the map $\rho:\mathcal{X}\to\mathcal{X}^{b}$ from elements in $\mathcal{X}$ to their canonical forms:
@@
In this example, we demonstrate solving for the moving frame for a system with both translation and rotation symmetries. Consider a car with state $x=(y,z,v_{y},v_{z},h_{y},h_{z})$ where $y$ and $z$ correspond to positions in a plane, $v_{y}$ and $v_{z}$ correspond to velocities along the $y$ and $z$ directions, and $h_{y}$ and $h_{z}$ correspond to the cosine and sine of the heading angle (and thus $h_{y}^{2}+h_{z}^{2}=1$). Since we assume the dynamics of the car are both translation- and rotation-invariant, we consider the special Euclidean group $G=\mathrm{SE}(2)$, and parameterize it by a translation $(y^{\prime},z^{\prime})$ and a rotation $\theta^{\prime}$. Then, with rotation matrix $R_{\theta^{\prime}}=\begin{bmatrix}\cos\theta^{\prime}&-\sin\theta^{\prime}\\ \sin\theta^{\prime}&\cos\theta^{\prime}\end{bmatrix}$, we define the action of $G$ on the state space by
$\phi_{(y^{\prime},z^{\prime},\theta^{\prime})}(x)=\begin{bmatrix}R_{\theta^{\prime}}&0&0\\ 0&R_{\theta^{\prime}}&0\\ 0&0&R_{\theta^{\prime}}\end{bmatrix}x+\begin{bmatrix}\begin{bmatrix}y^{\prime}\\ z^{\prime}\end{bmatrix}\\ 0\\ 0\end{bmatrix}.$ (3)
We now derive $\gamma$ and $\rho$. Let $\mathcal{X}^{a}$ be the projection of $\mathcal{X}$ onto the $(y,z,h_{y},h_{z})$ components. Pick $c=(0,0,1,0)$, representing a position at the origin with a heading angle of 00. There exists a unique $(y^{\prime},z^{\prime},\theta^{\prime})$ to transform each $x\in\mathcal{X}$ into the cross-section defined by $\mathcal{C}$. We solve the following system of equations to find the moving frame:
$\begin{bmatrix}0\\ 0\\ 1\\ 0\end{bmatrix}=\phi_{\gamma(x)}^{a}(x)=\begin{bmatrix}R_{\theta^{\prime}}&0\\ 0&R_{\theta^{\prime}}\end{bmatrix}\begin{bmatrix}y\\ z\\ h_{y}\\ h_{z}\end{bmatrix}+\begin{bmatrix}y^{\prime}\\ z^{\prime}\\ 0\\ 0\end{bmatrix}.$
@@
$\gamma(x)=\begin{bmatrix}-yh_{y}-zh_{z}\\ yh_{z}-zh_{y}\\ \mathrm{arctan2}(-h_{z},h_{y})\end{bmatrix}.$ (4)
@@ IV-A Two Cars Parking Scenario
The results of observation error vs. number of parameter updates are shown in Figure 4. At the lower number of parameters, the dynamical model training with symmetry outperforms the default method without symmetry by achieving a lower observation error. When there are two hidden layers, the method with symmetry learns faster than the default method without symmetry, although eventually the method without symmetry catches up in performance. When there are three hidden layers, the method without symmetry exceeds the performance of that of the method with symmetry. This suggests that the relative benefit of leveraging symmetry compared to the benefit from the number of parameters diminishes as the number of neural network parameters increases. One possible cause for this is that our method of leveraging symmetry removes redundant information, but the neural network might be able to exploit this redundant information when it has a sufficiently high number of parameters.

### v3

@@ Abstract
Recent work in reinforcement learning has leveraged symmetries in the model to improve sample efficiency in training a policy. A commonly used simplifying assumption is that the dynamics and reward both exhibit the same symmetry; however, in many real-world environments, the dynamical model exhibits symmetry independent of the reward model. In this paper, we assume only the dynamics exhibit symmetry, extending the scope of problems in reinforcement learning and learning in control theory to which symmetry techniques can be applied. We use Cartan’s moving frame method to introduce a technique for learning dynamics that, by construction, exhibit specified symmetries. Numerical experiments demonstrate that the proposed method learns a more accurate dynamical model.
@@ I Introduction
One idea gaining recent attention in RL methods is to integrate symmetries into model-based information. Using knowledge of symmetries, one can extrapolate observed behavior to other scenarios. Recent work has leveraged symmetries via equivariant networks [vanderpol2021mdp, wang_mathrmso2-equivariant_2022], which enforce certain transformation relationships on the inputs and outputs, and via data augmentation [weissenbacher2022koopman], where artificial data is created based on observed data and known transformations for training. However, a limitation of equivariant networks is that they must be designed for the particular group of symmetries to which they are equivariant. Unlike the above methods, which assume that the symmetries are known apriori, [weissenbacher2022koopman] learns symmetries from data and augments the training data using these symmetries.
@@
A common assumption when using symmetry in RL and optimal control is that both the dynamics and reward (or cost) function exhibit the same symmetries, which results in the optimal policy (or control law) likewise being symmetric [ravindran_model_2002, maidens_dynamic_programming]. The symmetric policy structure is exploited for Markov Decision Processes (MDPs) and RL in [vanderpol2021mdp, wang_mathrmso2-equivariant_2022, ravindran_model_2002, zinkevich_symmetry_2001], and for dynamic programming in [maidens_dynamic_programming]. In many scenarios, however, symmetries in the dynamical model are not necessarily applicable to the reward model. For example, multiple tasks might be executed in the same environment, where the same dynamical symmetries exist, while the reward model changes from task to task. Exploiting symmetries only in the dynamical model widens the scope of RL problems to which symmetry can be applied.
@@
This paper is organized as follows: in Section II, we provide background information on symmetries, and then use Cartan’s method of moving frames to relate a symmetrical dynamical model to a function operating on an input space of reduced dimension. In Section III, we use the result of Section II to learn a dynamical model that, by construction, satisfies specified symmetries. In Section IV we apply symmetry reduction to learn dynamical models for two examples.
@@ II Symmetries in Model Dynamics
We represent symmetries of (1) with an $r$-dimensional Lie group $G$ which is a Lie transformation group of both $\mathcal{X}$ and $\mathcal{U}$. Assume $r\leq n$. For notation, we consider the maps $\phi_{g}:\mathcal{X}\to\mathcal{X}$ defined by $x\mapsto g\cdot x$ and $\psi_{g}:\mathcal{U}\to\mathcal{U}$ defined by $u\mapsto g\cdot u$. Since the group actions are smooth, both of these maps are smooth for all $g\in G$. Further, they have smooth inverses $\phi^{-1}_{g}=\phi_{g^{-1}}$ and $\psi^{-1}_{g}=\psi_{g^{-1}}$.
@@ II-A Cartan’s Moving Frame
Consider the Lie group $G$ and its action on $\mathcal{X}$ as before. Assume $\mathcal{X}$ can be split into $\mathcal{X}^{a}$ and $\mathcal{X}^{b}$ with $r$ and $n-r$ dimensions respectively such that, for some $x_{0}\in\mathcal{X}$, the map from $g\in G$ to the projection of $g\cdot x_{0}$ onto $\mathcal{X}^{a}$ is invertible. Pick $c\in\mathcal{X}^{a}$ in the range of this map. For all $g\in G$, let $\phi_{g}^{a}:\mathcal{X}\to\mathcal{X}^{a}$ and $\phi_{g}^{b}:\mathcal{X}\to\mathcal{X}^{b}$ denote the compositions of $\phi_{g}$ and the projections onto $\mathcal{X}^{a}$ and $\mathcal{X}^{b}$ respectively. Assume that, for all $x\in\mathcal{X}$, there exists a unique $g\in G$ such that $\phi_{g}^{a}(x)=c$. The set $\mathcal{C}=\{x|\phi_{e}^{a}(x)=c\}$ is an $n-r$ dimensional submanifold of $\mathcal{X}$ called a cross-section. We can then define the unique map $\gamma:\mathcal{X}\to G$, such that $\phi_{\gamma(x)}(x)\in\mathcal{C}$. This $\gamma$ is called a moving frame. Using $\gamma$, we define the map $\rho:\mathcal{X}\to\mathcal{X}^{b}$ from elements in $\mathcal{X}$ to their canonical forms:
@@
In this example, we demonstrate solving for the moving frame for a system with both translation and rotation symmetries. Consider a car with state $x=(y,z,v_{y},v_{z},h_{y},h_{z})$ where $y$ and $z$ correspond to positions in a plane, $v_{y}$ and $v_{z}$ correspond to velocities along the $y$ and $z$ directions, and $h_{y}$ and $h_{z}$ correspond to the cosine and sine of the heading angle (and thus $h_{y}^{2}+h_{z}^{2}=1$). Since we assume the dynamics of the car are both translation- and rotation-invariant, we consider the special Euclidean group $G=\mathrm{SE}(2)$, and parameterize it by a translation $(\tilde{y},\tilde{z})$ and a rotation $\tilde{\theta}$. Then, with rotation matrix $R_{\tilde{\theta}}=\begin{bmatrix}\cos\tilde{\theta}&-\sin\tilde{\theta}\\ \sin\tilde{\theta}&\cos\tilde{\theta}\end{bmatrix}$, we define the action of $G$ on the state space by
$\phi_{(\tilde{y},\tilde{z},\tilde{\theta})}(x)=\begin{bmatrix}R_{\tilde{\theta}}&0&0\\ 0&R_{\tilde{\theta}}&0\\ 0&0&R_{\tilde{\theta}}\end{bmatrix}x+\begin{bmatrix}\begin{bmatrix}\tilde{y}\\ \tilde{z}\end{bmatrix}\\ 0\\ 0\end{bmatrix}.$ (3)
We now derive $\gamma$ and $\rho$. Let $\mathcal{X}^{a}$ be the projection of $\mathcal{X}$ onto the $(y,z,h_{y},h_{z})$ components. Pick $c=(0,0,1,0)$, representing a position at the origin with a heading angle of 00. There exists a unique $(\tilde{y},\tilde{z},\tilde{\theta})$ to transform each $x\in\mathcal{X}$ into the cross-section defined by $\mathcal{C}$. We solve the following system of equations to find the moving frame:
$\begin{bmatrix}0\\ 0\\ 1\\ 0\end{bmatrix}=\phi_{\gamma(x)}^{a}(x)=\begin{bmatrix}R_{\tilde{\theta}}&0\\ 0&R_{\tilde{\theta}}\end{bmatrix}\begin{bmatrix}y\\ z\\ h_{y}\\ h_{z}\end{bmatrix}+\begin{bmatrix}\tilde{y}\\ \tilde{z}\\ 0\\ 0\end{bmatrix}.$
@@
$\gamma(x)=\begin{bmatrix}\tilde{y}\\ \tilde{z}\\ \tilde{\theta}\end{bmatrix}=\begin{bmatrix}-yh_{y}-zh_{z}\\ yh_{z}-zh_{y}\\ \mathrm{arctan2}(-h_{z},h_{y})\end{bmatrix}.$ (4)
@@ IV-A Two Cars Parking Scenario
The results of observation error vs. number of parameter updates are shown in Figure 4. At the lower number of parameters, the dynamical model training with symmetry outperforms the default method without symmetry by achieving a lower observation error. When there are two hidden layers, the method with symmetry learns faster than the default method without symmetry, although eventually the method without symmetry catches up in performance. When there are three hidden layers, the method without symmetry exceeds the performance of the method with symmetry. This suggests that the relative benefit of leveraging symmetry compared to the benefit from the number of parameters diminishes as the number of neural network parameters increases. One possible cause for this is that our method of leveraging symmetry removes redundant information. This redundant information is a transformed version of the other available information, and might be more immediately useful to the network. As a result, the neural network does not have to learn these transformations on its own, making the redundant information beneficial when a sufficiently high number of parameters is available.